A clot blocks the right carotid artery, increasing the total resistance in the brain to 200 from 114 [mmHg.min.L-1 ] . What is the percent change in bloodflow to the brain?
Answer #1

F = P/R

F= blood flow

P= pressure difference

R= Resistance

R is decreased from 200 to 114

Initial flow = P/200

New flow = P/114

% change = P/114 - P/200 / P/200 * 100

= 86 P/22800 *20000/ P

= 75.43 %